23.28:At a certain distance from a point charge, the potential and electric field magnitude due to that charge are 4.98 V and 12.0 V/m, respectively. (Take the potential to be zero at infinity.)

1.What is the distance to the point charge? (d= ? m)
2.What is the magnitude of the charge? (q= ? c)
Answer #1
Concepts and reason

The concept required to solve this problem are electric field due to a point charge and electric potential due to a point charge.

Calculate distance d, solve the expression for both electric field and electric potential.

Fundamentals

Electric field is a vector field and at a point is the measure of force experienced by unit positive charge placed at that point by some source charge.

Electric field E at a point d distance apart from charge q1{q_1} is given as follows:

E=kq1d2r^\vec E = \frac{{k{q_1}}}{{{d^2}}}\hat r

Here, k is the proportionality factor and possess a constant value for a medium.

Electric potential at a point is the amount of work done in bringing a unit positive charge from a reference point generally at infinity to a specific point in the electric field without producing any acceleration.

Electric potential at a point d distance apart from charge q1{q_1} is given by following expression:

V=kq1dV = \frac{{k{q_1}}}{d}

The expression for source charge is given as follows:

q1=Vdk{q_1} = \frac{{Vd}}{k}

(a)

Consider a point at distance d from a point charge q1{q_1} .

The expression for magnitude of electric field E due to a point charge is given as follows:

E=kq1d2E = \frac{{k{q_1}}}{{{d^2}}}

The expression for electric potential is given by following expression:

V=kq1dV = \frac{{k{q_1}}}{d}

Substitute the expression for potential in formula of electric field as follows:

E=(kq1d)1d=V(1d)\begin{array}{c}\\E = \left( {\frac{{k{q_1}}}{d}} \right)\frac{1}{d}\\\\ = V\left( {\frac{1}{d}} \right)\\\end{array}

The relation between electric field and electric potential is,

E=VdE = \frac{V}{d}

Rearrange the equation for distance.

d=VEd = \frac{V}{E}

Substitute (12.0Vm1)\left( {12.0{\rm{ V}} \cdot {{\rm{m}}^{ - 1}}} \right) for E and (4.98V)\left( {{\rm{4}}{\rm{.98 V}}} \right) for V as follows:

d=VE=(4.98V)(12.0Vm1)=0.415m\begin{array}{c}\\d = \frac{V}{E}\\\\ = \frac{{\left( {{\rm{4}}{\rm{.98 V}}} \right)}}{{\left( {12.0{\rm{ V}} \cdot {{\rm{m}}^{ - 1}}} \right)}}\\\\ = 0.415{\rm{ m}}\\\end{array}

(b)

Find the magnitude of the source charge.

The expression for electric potential is given by following expression:

V=kq1dV = \frac{{k{q_1}}}{d}

Rearrange the expression for potential for the charge as follows:

q1=Vdk{q_1} = \frac{{Vd}}{k}

Substitute (9×109Nm2C2)\left( {9 \times {{10}^9}{\rm{ N}} \cdot {{\rm{m}}^2} \cdot {{\rm{C}}^{ - 2}}} \right) for k, (0.415m)\left( {0.415{\rm{ m}}} \right) for d and (4.98V)\left( {{\rm{4}}{\rm{.98 V}}} \right) for V as follows:

q1=(4.98V)(0.415m)(9×109Nm2C2)=0.23×109C\begin{array}{c}\\{q_1} = \frac{{\left( {{\rm{4}}{\rm{.98 V}}} \right)\left( {0.415{\rm{ m}}} \right)}}{{\left( {9 \times {{10}^9}{\rm{ N}} \cdot {{\rm{m}}^2} \cdot {{\rm{C}}^{ - 2}}} \right)}}\\\\ = 0.23 \times {10^{ - 9}}{\rm{ C}}\\\end{array}

Ans: Part a

The distance of a point charge from field point is (0.415m)\left( {0.415{\rm{ m}}} \right) .

Part b

The magnitude of source charge q1{q_1} is (0.23×109C)\left( {0.23 \times {{10}^{ - 9}}{\rm{ C}}} \right) .